OpenStudy (anonymous):

Carson eats 2 1/4 ounce of cheese. One serving of cheese is 1/3 ounce. How many servings did Carson eat?

OpenStudy (anonymous):

@ghazi

OpenStudy (bahrom7893):

how many 1/3s go into 2 1/4?

OpenStudy (bahrom7893):

2 1/4 = (4*2+1)/4 = 9/4

OpenStudy (anonymous):

2.25 pounds?

OpenStudy (bahrom7893):

(9/4)/(3/4)

OpenStudy (anonymous):

ounce*

OpenStudy (bahrom7893):

he had 3 servings

OpenStudy (anonymous):

yeah. k.

OpenStudy (anonymous):

just realized that.

OpenStudy (anonymous):

Thanks.